Tahir measured the distance around a square field as 200 rods (lathi). Later he found that the length of this rod was 140cm. Find the side of this field in metres.

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

From the question it is given that,

Tahir measured the distance around a square field as 200 rods Length of the rod = 140 cm
So, total distance of square field = 200 × 140

= 28000 cm
Then, length of one side of square field = 28000/4

= 7000 cm
We know that, 1 m = 100 cm Therefore, side of field = 7000/100

= 70 m
Thanks